# Detecting Invariant Manifolds in ReLU-Based RNNs

### v1

@@ 1 Introduction
Recurrent neural networks (RNNs) are widely employed for time series forecasting (Park et al., 2018; Gu and Dao, 2023) and dynamical systems (DS) reconstruction (Hess et al., 2023; Brenner et al., 2022, 2024b; Platt et al., 2023), especially in scientific applications like climate modeling (Patel and Ott, 2023) or neuroscience (Durstewitz et al., 2023), as well as in medical domains. RNNs experienced a recent revival due to the advance of new powerful training algorithms that avoid vanishing or exploding gradients (Hess et al., 2023), and novel network architectures (Schmidt et al., 2021; Rusch and Mishra, 2020; Peng et al., 2023; Gu and Dao, 2023; Rusch et al., 2022), in particular in the context of state space models (which are essentially linear RNNs with nonlinear readouts and input gating; (Gu and Dao, 2023; Orvieto et al., 2023)). What lags behind, like in many other areas of deep learning, is a thorough theoretical understanding of the behavior of these systems and how they achieve the tasks they were trained on. Yet, such an understanding is crucial especially in scientific and medical areas where we are often interested in using trained RNNs as surrogate models for the underlying DS, providing mechanistic insight into the dynamical processes that generated the observed time series.
@@
Formally, RNNs are recursive maps, hence discrete-time dynamical systems (Mikhaeil et al., 2022). Dynamical systems theory (DST) offers a rich repertoire of mathematical tools for analyzing the behavior of such systems (Guckenheimer and Holmes, 2013). Yet, the exploration of high-dimensional RNN dynamics remains a challenge due to limitations in current numerical methods, which struggle to scale and tend to yield only approximate results (Katz and Reggia, 2018; Golub and Sussillo, 2018).
@@
The dynamical behavior of a system is governed by its state space topology and geometry (Hasselblatt and Katok, 2002), most prominently topological objects like attractors, such as stable fixed points, cycles, or chaotic sets, which determine the system’s long-term behavior. Similarly important are the stable and unstable manifolds of fixed points and cycles, although they received much less attention in the scientific and ML communities. Stable manifolds are the sets of points in a system’s state space that converge towards an equilibrium or periodic orbit in forward-time, while unstable manifolds, conversely, are the sets of points that converge to an equilibrium or periodic orbit in backward-time. Stable manifolds of saddle points delineate the boundaries between basins of attraction in multistable systems, that is, systems that harbor multiple attractor objects between which it can be driven back and forth by perturbations like external inputs or noise (Feudel et al., 2018). Multistability has been hypothesized to be an important computational property of RNNs, most prominently in computational neuroscience where it has been linked to working memory (Durstewitz et al., 2000) or decision making (Wang, 2008). Different attractors may, for instance, correspond to different active memory states, and the process of convergence to one of these attractors to memory retrieval and pattern completion, as in Hopfield networks (Hopfield, 1982).
@@
Tracing out stable and unstable manifolds is also important for finding homo- and heteroclinic orbits, which connect a cyclic point to itself or to another such point, respectively. These orbits provide a skeleton for the dynamics, forming structures like separatrix cycles (Perko, 2001) and heteroclinic channels (Rabinovich et al., 2008) which have been implied in flexible sequence generation. Intersections between stable and unstable manifolds also give rise to so-called homoclinic points which create sensitive regions in a system’s state space associated with chaos (Wiggins, 1988). Chaotic behavior, in turn, or regimes at the edge-of-chaos, have been associated with increased expressivity (larger function classes that can be emulated) and computational power in RNNs (Siegelmann and Sontag, 1992; Bertschinger and Natschläger, 2004; Pereira-Obilinovic et al., 2023). Chaos is also a practical problem in training RNNs, since it causes exploding gradients if not taken care of by special training techniques (Mikhaeil et al., 2022; Hess et al., 2023).
@@
Here we introduce a novel algorithm for computing the stable and unstable manifolds of fixed and cyclic points for the class of piecewise-linear (PL) RNNs (PLRNNs), which use piecewise nonlinearities like the rectified-linear unit (ReLU) as their activation function. Unlike traditional techniques designed for smooth dynamical systems—such as numerical continuation methods (Krauskopf et al., 2007) —the proposed method explicitly exploits the piecewise-linear structure of ReLU-based PLRNNs to enable the exact location of stable and unstable manifolds. In contrast to smooth systems, methods for discrete-time systems are much more scarce, especially if these involve discontinuities in their Jacobians like ReLU-based RNNs. PLRNNs have emerged as one of the most powerful classes of models for dynamical systems reconstruction (Hess et al., 2023; Brenner et al., 2022, 2024b; Nassar et al., 2019), partly because the linear subspaces of such models support the indefinite online retention of memory contents without exploding or vanishing gradients (Schmidt et al., 2021; Orvieto et al., 2023; Gu and Dao, 2023), and partly because their PL structure makes them more tractable (Coombes et al., 2024). For instance, efficient algorithms for exactly localizing fixed and cyclic points in polynomial time exist (Eisenmann et al., 2023), on which we will partly build here. We illustrate how the algorithm can be employed to delineate the boundaries of basins of attraction and to detect homoclinic intersections, thus establishing the existence of chaos, and show how it can be used on empirical data to gain insight into dynamical mechanisms.
@@ 2 Related work
While there is an extensive literature by now on dynamical systems reconstruction with RNNs (Brenner et al., 2022, 2024a; Hess et al., 2023; Patel and Ott, 2023; Platt et al., 2023; Cestnik and Abel, 2019), work on algorithms for dissecting topological properties of trained systems is much more scarce. Existing research is almost exclusively focused on algorithms for locating stable and unstable fixed points and cycles (Golub and Sussillo, 2018; Eisenmann et al., 2023), but not the invariant manifolds associated with them, despite their crucial role in structuring the state space and dynamics. More generally in the DST literature, a related class of methods are so-called continuation methods which numerically trace back important curves and manifolds in dynamical systems, mostly for systems of ordinary differential equations (Krauskopf et al., 2007; Osinga, 2014). Recently methods have been developed specifically for PL maps (Simpson, 2023), but, similar to continuation methods, they seriously suffer from the curse of dimensionality. Thus, current methods effectively work only for very low-dimensional ($\leq 5d$) systems, in contrast to the size of many modern RNNs used for time series forecasting or dynamical systems reconstruction. The specific structure of PL maps, of which all types of PLRNNs are specific examples, considerably eases the computation of certain topological properties. This also provided the basis for the SCYFI algorithm for locating fixed points and cycles, which scales polynomially, in fact often linearly, with the dimensionality of the PLRNN’s latent space (Eisenmann et al., 2023). To our knowledge, no existing method efficiently detects stable and unstable manifolds in discrete-time RNNs. Our approach fills this gap by leveraging the piecewise-linear nature of ReLU-based RNNs to construct manifolds directly, bypassing the limitations of traditional techniques.
@@ 3.1 PLRNN architectures
A PLRNN (Durstewitz, 2017) is simply a ReLU-based RNN with an additional linear term, which takes the basic form
@@
where ${\bm{A}}\in\mathbb{R}^{M\times M}$ is a diagonal matrix, ${\bm{W}}\in\mathbb{R}^{M\times M}$ an off-diagonal matrix, the element-wise ReLU non-linearity $\Phi(\cdot)=\text{max}(0,\cdot)$, and ${\bm{h}}\in\mathbb{R}^{M}$ a bias term. Several variants of the PLRNN have been introduced to enhance its expressivity or reduce its dimensionality (Brenner et al., 2022), of which we picked for demonstration here the shallow PLRNN (shPLRNN; (Hess et al., 2023)), ${\bm{z}}_{t}={\bm{A}}{\bm{z}}_{t-1}+{\bm{W}}_{1}\Phi({\bm{W}}_{2}{\bm{z}}_{t-1}+{\bm{h}}_{2})+{\bm{h}}_{1}$ with ${\bm{W}}_{1}\in\mathbb{R}^{M\times H}$, ${\bm{W}}_{2}\in\mathbb{R}^{H\times M}$, and the recently proposed almost-linear RNN (ALRNN; (Brenner et al., 2024a)), which attempts to use as few nonlinearities as needed for the problem at hand, $\Phi({\bm{z}}_{t})=\bqty{{\bm{z}}_{1,t},\ldots,{\bm{z}}_{M-P,t},\text{max}(0,{\bm{z}}_{M-P+1,t}),\ldots,\text{max}(0,{\bm{z}}_{M,t})}$.
The piecewise linear structure of these PLRNNs can be exposed by rewriting the ReLU in Equation˜1 as
@@ 3.2 Mathematical preliminaries
Let $F:\mathbb{R}^{M}\to\mathbb{R}^{M}$ be a map and $\bm{p}$} be a hyperbolic period-$m$ cyclic point of $F$. The local stable manifold of $\bm{p}$}, $W_{loc}^{s}({\bm{p}})$, is defined as
$W_{loc}^{s}({\bm{p}})\,:=\,\{{\bm{x}}\in\mathbb{R}^{M}\,:\,F^{nm}({\bm{x}})\to{\bm{p}}\quad\text{as}\quad n\to\infty\}.$
The local unstable manifold of $\bm{p}$}, $W_{loc}^{u}({\bm{p}})$, is defined as
$W_{loc}^{u}({\bm{p}})\,:=\,\{{\bm{x}}\in\mathbb{R}^{M}\,:\,F^{-nm}({\bm{x}})\to{\bm{p}}\quad\text{as}\quad n\to\infty\}.$
The global stable manifold is the union of all preimages of the local stable manifold, and the global unstable manifold is created by the union of all images (forward iterations) of the local unstable manifold (Patra and Banerjee, 2018):
$W^{s}({\bm{p}})\,:=\,\bigcup_{n=1}^{\infty}F^{-n}\Big(W_{loc}^{s}({\bm{p}})\Big),\hskip 22.76228ptW^{u}({\bm{p}})\,:=\,\bigcup_{n=1}^{\infty}F^{n}\Big(W_{loc}^{u}({\bm{p}})\Big).$ (3)
@@
In each locally linear region of a PL dynamical system, the state space decomposes into invariant subspaces for the sets of positive and negative eigenvalues, respectively. For real eigenvalues, trajectories lie on hyperplanes on which the motion is directed towards or away from the fixed point in a straight line. More generally, complex eigenvectors can occur and cause curvature, or eigenvalues may repeat, and the general dynamics in each affine subregion (${\bm{z}}_{t}=\tilde{{\bm{W}}}{\bm{z}}_{t-1}+{\bm{h}},\tilde{{\bm{W}}}:={\bm{A}}+{\bm{W}}{\bm{D}}({\bm{z}}_{t-1})$) is given by
${\bm{z}}_{t}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}c_{j}\lambda_{j}^{t}{\bm{v}}_{j}\;+\;\lambda_{i}^{t}\sum_{r=1}^{m}c_{r}\frac{t^{r-1}}{(r-1)!}{\bm{w}}_{r}\;+\;\sum_{k=0}^{t-1}\tilde{{\bm{W}}}^{k}{\bm{h}}.$ (4)
where the $c_{i}\in\mathbb{R}$ are determined from the initial condition, ${\bm{v}}_{i}$ are the eigenvectors, ${\bm{w}}_{i}$ are generalized eigenvectors (associated with eigenvalues with geometric multiplicity less than the algebraic, and for multiplicity of $m>1$). This formula describes the exponential evolution for all non-defective eigenvalues $\lambda_{j}$ and polynomially modulated exponentials for the defective eigenvalue $\lambda_{i}$, see Appx. C for details and special cases (Perko, 2001; Hirsch et al., 2013).
Stable manifolds of saddle objects segregate the state space into different basins of attraction (Ganatra and Banerjee, 2022), which are sets of points from which the state evolves toward one or the other attractor. Formally, they are given by (Alligood et al., 1996):
@@
Stable and unstable manifolds are crucially important for the system dynamics not only because they dissect the state space into different regions of flow, but also because their intersections can give rise to complex types of dynamics like heteroclinic channels or chaos (Rabinovich et al., 2008; Perko, 2001).
@@
$\mathcal{O}_{hom}\subset W^{s}({\bm{p}})\;\cap\;W^{u}({\bm{p}}).$
If ${\bm{x}}\neq{\bm{p}}$ is a point where the stable and unstable manifolds of $\bm{p}$} intersect, then $\bm{x}$} is referred to as a homoclinic point or intersection (Patra and Banerjee, 2018; Perko, 2001). Such an intersection of the stable and unstable manifolds leads to a horseshoe structure associated with a fractal geometry and chaos (see Appx. Def. 5, (Wiggins, 1988)).
@@
Let $\bm{p}$} and $\bm{q}$} be two distinct saddle fixed points (or saddle cycles) of the map $F$. A heteroclinic orbit $\mathcal{O}_{het}=\{{\bm{x}}_{n}\}_{n\in\mathbb{N}}$ connects two different fixed points $\bm{p}$} and $\bm{q}$}, that is, ${\bm{x}}_{n}\to{\bm{p}}$ as $n\to+\infty$ and ${\bm{x}}_{n}\to{\bm{q}}$ as $n\to$-$\infty$. Heteroclinic intersections (or points) can be defined similarly to homoclinic intersections, and, like homoclinic points, inevitably lead to chaos (Wiggins, 1988; Patra and Banerjee, 2018).

### v3

@@ 1 Introduction
Recurrent neural networks (RNNs) are widely employed for time series forecasting (Park; gu2023mamba) and dynamical systems (DS) reconstruction (hess; brenner_tractable_2022; brennerintegrating; platt2023constraining), especially in scientific applications like climate modeling (patel) or neuroscience (durstewitz2023reconstructing), as well as in medical domains. RNNs experienced a recent revival due to the advance of new powerful training algorithms that avoid vanishing or exploding gradients (hess), and novel network architectures (sch; rusch2020coupled; peng2023rwkv; gu2023mamba; ruschlong), in particular in the context of state space models (which are essentially linear RNNs with nonlinear readouts and input gating; (gu2023mamba; orvieto2023resurrecting)). What lags behind, like in many other areas of deep learning, is a thorough theoretical understanding of the behavior of these systems and how they achieve the tasks they were trained on. Yet, such an understanding is crucial especially in scientific and medical areas where we are often interested in using trained RNNs as surrogate models for the underlying DS, providing mechanistic insight into the dynamical processes that generated the observed time series.
@@
Formally, RNNs are recursive maps, hence discrete-time dynamical systems (mikhaeil2021difficulty). Dynamical systems theory (DST) offers a rich repertoire of mathematical tools for analyzing the behavior of such systems (guckenheimer2013nonlinear). Yet, the exploration of high-dimensional RNN dynamics remains a challenge due to limitations in current numerical methods, which struggle to scale and tend to yield only approximate results (katz_using_2018; golub_fixedpointfinder_2018).
@@
The dynamical behavior of a system is governed by its state space topology and geometry (hasselblatt2002handbook), most prominently topological objects like attractors, such as stable fixed points, cycles, or chaotic sets, which determine the system’s long-term behavior. Similarly important are the stable and unstable manifolds of fixed points and cycles, although they received much less attention in the scientific and ML communities. Stable manifolds are the sets of points in a system’s state space that converge towards an equilibrium or periodic orbit in forward-time, while unstable manifolds, conversely, are the sets of points that converge to it in backward-time. Stable manifolds of saddle points delineate the boundaries between basins of attraction in multistable systems, that is, systems that harbor multiple attractor objects between which it can be driven back and forth by perturbations like external inputs or noise (feu18). Multistability has been hypothesized to be an important computational property of RNNs, most prominently in computational neuroscience where it has been linked to working memory (dur0) or decision making (wang2008decision).
@@
Tracing out stable and unstable manifolds is also important for finding homo- and heteroclinic orbits, which connect a cyclic point to itself or to another such point, respectively. These orbits provide a skeleton for the dynamics, forming structures like separatrix cycles (per) and heteroclinic channels (rabinovich2008transient) which have been implied in flexible sequence generation. Intersections between stable and unstable manifolds further give rise to so-called homoclinic points which create sensitive regions in a system’s state space associated with chaos (wig). Chaotic behavior, in turn, or regimes at the edge-of-chaos, have been associated with increased expressivity (larger function classes that can be emulated) and computational power in RNNs (siegelmann1992computational; bertschinger2004real; Pereira) (mikhaeil2021difficulty; hess).
@@
Here we introduce a novel algorithm for computing the stable and unstable manifolds of fixed and cyclic points for the class of piecewise-linear (PL) RNNs (PLRNNs), which use piecewise nonlinearities like the ReLU as their activation function. Unlike traditional techniques designed for smooth dynamical systems—such as numerical continuation methods (krauskopf2007numerical) —the proposed method explicitly exploits the piecewise-linear structure of PLRNNs to enable the exact location of stable and unstable manifolds. In contrast to ODE systems, methods for discrete-time systems are much more scarce, especially if these involve discontinuities in their Jacobians like ReLU-based RNNs. PLRNNs have emerged as one of the most powerful classes of models for DS reconstruction (hess; brenner_tractable_2022; brennerintegrating; nassartree), partly because the linear subspaces of such models support the indefinite online retention of memory contents without exploding or vanishing gradients (sch; orvieto2023resurrecting; gu2023mamba), and partly because their PL structure makes them more tractable (coombes2024osc).In fact, PL systems have been popular in engineering (bemporad2000piecewise; carmona2002simplifying; storace2004piecewise) and mathematics (alligood1996; avrutin_continuous_2019; coombes2024osc; simpson2023compute) for decades for these reasons, and a variety of PL models similar in structure to PLRNNs exist, like threshold-linear networks (TLNs; curto2023graph; parmelee2022core; morrison2016diversity; hahnloser2000permitted), switching linear DS (SLDS; linderman_recurrent_2016; linderman_bayesian_2017), or Lur’e systems in control theory (su2023necessity; waitman2017incremental; bill2016stability), with a substantial body of mathematical theory built around them (parmelee2022core; morrison2016diversity; brenner2024alrnn; avrutin_continuous_2019; amaral_piecewise_2006; simpson2023compute; coombes2024osc; casdagli1989nonlinear; ives2012detecting; costa2019adaptive). Further, efficient algorithms for exactly localizing fixed and cyclic points in polynomial time exist (eisenmann_bifurcations_2023), on which we will build here. We illustrate how our algorithm can be employed to delineate the boundaries of basins of attraction and to detect homoclinic intersections, thus establishing the existence of chaos, and show how it can be used on empirical data to gain insight into dynamical mechanisms.
@@ 2 Related work
While there is an extensive literature by now on DS reconstruction with RNNs (brenner_tractable_2022; brenner2024alrnn; hess; patel; platt2023constraining; Cestnik), work on algorithms for dissecting topological properties of trained systems is much more scarce. Existing research is almost exclusively focused on algorithms for locating stable and unstable fixed points and cycles (golub_fixedpointfinder_2018; eisenmann_bifurcations_2023), but not the invariant manifolds associated with them, despite their crucial role in structuring the state space and dynamics. More generally in the DST literature, a related class of methods are so-called continuation methods which numerically trace back important curves and manifolds in dynamical systems, mostly for systems of ordinary differential equations (krauskopf2007numerical; osinga2014computing). Recently methods have been developed specifically for PL maps (simpson2023compute), but, similar to continuation methods, they seriously suffer from the curse of dimensionality. Thus, current methods effectively work only for very low-dimensional ($\leq 5d$) systems, in contrast to the size of many modern RNNs used for time series forecasting or dynamical systems reconstruction. The specific structure of PL maps, of which all types of PLRNNs are specific examples, considerably eases the computation of certain topological properties. This also provided the basis for the SCYFI algorithm for locating fixed points and cycles, which scales polynomially, in fact often linearly, with the dimensionality of the PLRNN’s latent space (eisenmann_bifurcations_2023). To our knowledge, no existing method efficiently detects stable and unstable manifolds in discrete-time RNNs. Our approach fills this gap by leveraging the PL nature of ReLU-based RNNs to construct manifolds directly, bypassing the limitations of traditional techniques.
@@ 3.1 PLRNN architectures
A PLRNN (dur) is simply a ReLU-based RNN
@@
where ${\bm{A}}\in\mathbb{R}^{M\times M}$ is a diagonal matrix, ${\bm{W}}\in\mathbb{R}^{M\times M}$ an off-diagonal matrix, the element-wise ReLU non-linearity $\Phi(\cdot)=\text{max}(0,\cdot)$, and ${\bm{h}}\in\mathbb{R}^{M}$ a bias term. Several variants of the PLRNN have been introduced to enhance its expressivity or reduce its dimensionality (brenner_tractable_2022), of which we picked for demonstration here the shallow PLRNN (shPLRNN; (hess)), ${\bm{z}}_{t}={\bm{A}}{\bm{z}}_{t-1}+{\bm{W}}_{1}\Phi({\bm{W}}_{2}{\bm{z}}_{t-1}+{\bm{h}}_{2})+{\bm{h}}_{1}$ with ${\bm{W}}_{1}\in\mathbb{R}^{M\times H}$, ${\bm{W}}_{2}\in\mathbb{R}^{H\times M}$, and the recently proposed almost-linear RNN (ALRNN; (brenner2024alrnn)), which attempts to use as few nonlinearities as needed for the problem at hand, $\Phi({\bm{z}}_{t})=\bqty{{\bm{z}}_{1,t},\ldots,{\bm{z}}_{M-P,t},\text{max}(0,{\bm{z}}_{M-P+1,t}),\ldots,\text{max}(0,{\bm{z}}_{M,t})}$.
The PL structure of these PLRNNs can be exposed by rewriting the ReLU in Equation˜1 as
@@ 3.2 Mathematical preliminaries
Let $F:\mathbb{R}^{M}\to\mathbb{R}^{M}$ be a map and $\bm{p}$} be a hyperbolic period-$m$ cyclic point of $F$. The local stable manifold of $\bm{p}$}, $W_{loc}^{\{+1\}}({\bm{p}})$, is defined as
$W_{loc}^{\{+1\}}({\bm{p}})\,:=\,\{{\bm{x}}\in\mathbb{R}^{M}\,:\,F^{nm}({\bm{x}})\to{\bm{p}}\quad\text{as}\quad n\to\infty\}.$
The local unstable manifold of $\bm{p}$}, $W_{loc}^{\{-1\}}({\bm{p}})$, is defined as
$W_{loc}^{\{-1\}}({\bm{p}})\,:=\,\{{\bm{x}}\in\mathbb{R}^{M}\,:\,F^{-nm}({\bm{x}})\to{\bm{p}}\quad\text{as}\quad n\to\infty\}.$
The global stable manifold is the union of all preimages of the local stable manifold, and the global unstable manifold is created by the union of all images (forward iterations) of the local unstable manifold (pat2):
$W^{\{+1\}}({\bm{p}})\,:=\,\bigcup_{n=1}^{\infty}F^{-n}\Big(W_{loc}^{\{+1\}}({\bm{p}})\Big),\hskip 22.76228ptW^{\{-1\}}({\bm{p}})\,:=\,\bigcup_{n=1}^{\infty}F^{n}\Big(W_{loc}^{\{-1\}}({\bm{p}})\Big).$ (3)
@@
For a PLRNN, since it is a PL map, in each linear subregion we have the following explicit (non-recursive) expression for the dynamics along trajectories (${\bm{z}}_{t}=\tilde{{\bm{W}}}{\bm{z}}_{t-1}+{\bm{h}},\tilde{{\bm{W}}}:={\bm{A}}+{\bm{W}}{\bm{D}}({\bm{z}}_{t-1})$):
${\bm{z}}_{t}$ $=\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n}c_{j}\lambda_{j}^{t}{\bm{v}}_{j}\;+\;\lambda_{i}^{t}\sum_{r=1}^{d}c_{r}\frac{t^{r-1}}{(r-1)!}{\bm{w}}_{r}\;+\;\sum_{k=0}^{t-1}\tilde{{\bm{W}}}^{k}{\bm{h}}.$ (4)
where the $c_{i}\in\mathbb{R}$ are determined from the initial condition, ${\bm{v}}_{i}$ are the eigenvectors, ${\bm{w}}_{i}$ are generalized eigenvectors (associated with defective eigenvalues with geometric multiplicity less than the algebraic). This formula describes the exponential evolution for all non-defective eigenvalues $\lambda_{j}$ and polynomially modulated exponentials for the defective eigenvalue $\lambda_{i}$, see Appx. B for details (per; hirsch2013differential). While this expression is for single orbits, it describes how curvature in the manifolds can arise as we cross the boundaries of linear subregions (cf. sect. 3.3).
Stable manifolds of saddle objects segregate the state space into different basins of attraction (ganatra2022sketching), which are sets of points from which the state evolves toward one or the other attractor. Formally, they are given by (alligood1996):
@@
Stable and unstable manifolds are crucially important for the system dynamics not only because they dissect the state space into different regions of flow, but also because their intersections can give rise to complex types of dynamics like heteroclinic channels or chaos (rabinovich2008transient; per).
@@
$\mathcal{O}_{hom}\subset W^{\{+1\}}({\bm{p}})\;\cap\;W^{\{-1\}}({\bm{p}}).$
If ${\bm{x}}\neq{\bm{p}}$ is a point where the stable and unstable manifolds of $\bm{p}$} intersect, then $\bm{x}$} is referred to as a homoclinic point or intersection (pat2; per). Such an intersection of the stable and unstable manifolds leads to a horseshoe structure associated with a fractal geometry and chaos (see Appx. I, wig).
@@
Let $\bm{p}$} and $\bm{q}$} be two distinct saddle fixed points (or saddle cycles) of the map $F$. A heteroclinic orbit $\mathcal{O}_{het}=\{{\bm{x}}_{n}\}_{n\in\mathbb{N}}$ connects two different fixed points $\bm{p}$} and $\bm{q}$}, that is, ${\bm{x}}_{n}\to{\bm{p}}$ as $n\to+\infty$ and ${\bm{x}}_{n}\to{\bm{q}}$ as $n\to$-$\infty$. Heteroclinic intersections (or points) can be defined similarly to homoclinic intersections, and, like homoclinic points, inevitably lead to chaos (wig; pat2).